a 48 cm wire is to be cut into two pieces. One piece must be 10 cm longer than the shorter piece. How long should each piece be?
Answer by cheesey(22)

x = shorter piece
x + 10 = longer piece
x+(x+10)=48
2x + 10 = 48
2x = 38
x = 19
shorter piece is 19cm, longer piece is 19 + 10 = 29 cm.